OpenStudy (anonymous):

how do i solve (1-i)^6

OpenStudy (anonymous):

ıs i complex number?

OpenStudy (anonymous):

dude ,dont take the pervious answer it was wrong ,ma small brother fault !!!

OpenStudy (experimentx):

first change it into e^itheta form

OpenStudy (experimentx):

|1-i| <--- first find the mod => \( \sqrt{1^+(-1)^2} = \sqrt2\) => \( \sqrt2 ( \frac{1}{\sqrt 2} - \frac{1}{\sqrt 2} i )\) => \( \sqrt2 ( \cos (-\pi/4) + \sin (-\pi/4) i)\) => \( \sqrt2 ( e^{-i\pi/4})\) Now it should be easy.

OpenStudy (anonymous):

The answer I got was 8i after inserting your equation into a calculator.

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=%281-i%29%5E6